OpenStudy (anonymous):

Solve using substitution: 6y − x − 9 = 0 -4x = -5y − 2 A. No solution B. Infinitely many solutions C. (-4, -1/6) D. (3, 2)

OpenStudy (anonymous):

first, solve for "Y" in your second eqaution: -4x = -5y -2 -4(0) + 5y = -2 5y = -2 y = -2/5 now plug "Y" into your first equation and solve for "X" 6(-2/5) - x = 9 -12/5 - x = 9 -x = 33/5 x = -33/5 either i royally screwed up or there is no solution to this answer o.O (its probably the first one) you may want to clarify with someone

OpenStudy (anonymous):

@jollysailorbold ^^^

OpenStudy (anonymous):

hi!

OpenStudy (anonymous):

hello (: can you help me with this problem ?

OpenStudy (anonymous):

yeah give me a bit:)

OpenStudy (anonymous):

he royally screwed up because i got this: x=3 y=2

OpenStudy (anonymous):

unless i royally screwed up too.

OpenStudy (anonymous):

seems a lot better than mine at this point lol, i would go with jollies

OpenStudy (anonymous):

(-x+6y=9)4 -4x+5y=-2 this gives........... -4x+24y=36 -4x+5y=-2 on solving these two equations we get:- 19y=38 =>y=38/19 =>y=2 on putting y=2 in equation no 1, we get...... -x+6y=9 -x+6.2=9 =>x=3

OpenStudy (anonymous):

is this fine?????